# Dunkl approach to slice regular functions

### v3

@@ 4. Fueter’s theorem revisited
$\begin{split}\mathcal{V}_{n}f(x_{0}+\underline{x})\sim\left\{\begin{array}[]{ll}\Delta_{\mathbb{R}^{m+1}}^{n}f,&0\leq n\leq(m-1)/2\in\mathbb{N},\\ \Delta_{\kappa}^{n}f,&0\leq n\leq\gamma_{\kappa}+(m-1)/2\in\mathbb{N}.\\ \end{array}\right.\end{split}$
@@
(4.3) $\begin{split}\mathcal{V}_{n}f(x_{0}+\underline{x})\in\left\{\begin{array}[]{ll}\ker\overline{\partial}_{x},&n=(m-1)/2\in\mathbb{N},\\ \ker D_{\kappa},&n=\gamma_{\kappa}+(m-1)/2\in\mathbb{N}.\\ \end{array}\right.\end{split}$
@@
$\begin{matrix}&f\longrightarrow&\Delta_{\mathbb{R}^{m+1}}f\longrightarrow&\cdots&\longrightarrow\Delta_{\mathbb{R}^{m+1}}^{\frac{m-1}{2}}f&\longrightarrow\mathcal{V}_{\frac{m+1}{2}}f\longrightarrow&\mathcal{V}_{\frac{m+3}{2}}f\longrightarrow&\cdots\\ \\ \gamma_{\kappa}:&\frac{1-m}{2}\quad\quad&\frac{3-m}{2}\qquad&\cdots&\qquad 0&1&2&\cdots\\ \end{matrix}$
@@ 5. New approach to generate monogenic functions by the inverse intertwining operator
Suppose $f$ is a slice regular function of the form (4.2), then the function $\mathcal{T}f(x_{0}+\underline{x})$ is monogenic, i.e. it is in the kernel of the generalized Cauchy-Riemann operator ¯$\overline{\partial}_{x}=\partial_{x_{0}}+\sum_{j=1}^{m}e_{j}\partial_{x_{j}}$.
@@
Furthermore, if we consider the root system ={±$\mathscr{R}=\{\pm e_{j}\}_{j=1}^{3}$, with $D_{h}=\partial_{x_{0}}+e_{1}T_{1}+e_{2}T_{2}+e_{3}T_{3}$, in which
@@
¯$\overline{\partial}_{x}\mathcal{T}(x_{0}+\underline{x})^{2}=-4x_{0}\left(1+\sum_{i=1}^{3}\kappa(e_{i})\right)=0.$
@@
and the root system ={±$\mathscr{R}=\{\pm e_{i}\}_{i=1}^{m}$. Then we have
@@
¯$\overline{\partial}_{x}\mathcal{T}x^{2}=4x_{0}\left(\frac{1-m}{2}-\sum_{i=1}^{m}\kappa(e_{i})\right)=0.$

### v4

@@ 4. Fueter’s theorem revisited
[…]
@@
[…]
@@
$\begin{matrix}&f\longrightarrow&\Delta_{\mathbb{R}^{m+1}}f\longrightarrow&\cdots&\longrightarrow\Delta_{\mathbb{R}^{m+1}}^{\frac{m-1}{2}}f&\longrightarrow\mathcal{V}_{\frac{m+1}{2}}f\longrightarrow&\mathcal{V}_{\frac{m+3}{2}}f\longrightarrow&\cdots\\ \\ \gamma_{\kappa}:&\frac{1-m}{2}\quad\quad&\frac{3-m}{2}\qquad&\cdots&\qquad 0&1&2$$\cdots\\ \end{matrix}$
@@ 5. New approach to generate monogenic functions by the inverse intertwining operator
Suppose $f$ is a slice regular function of the form (4.2), then the function $\mathcal{T}f(x_{0}+\underline{x})$ is monogenic, i.e. it is in the kernel of the generalized Cauchy-Riemann operator $\overline{\partial}_{x}=\partial_{x_{0}}+\sum_{j=1}^{m}e_{j}\partial_{x_{j}}$.
@@
Furthermore, if we consider the root system $\mathscr{R}=\{\pm e_{j}\}_{j=1}^{3}$, with $D_{h}=\partial_{x_{0}}+e_{1}T_{1}+e_{2}T_{2}+e_{3}T_{3}$, in which
@@
$\overline{\partial}_{x}\mathcal{T}(x_{0}+\underline{x})^{2}=-4x_{0}\left(1+\sum_{i=1}^{3}\kappa(e_{i})\right)=0.$
@@
and the root system $\mathscr{R}=\{\pm e_{i}\}_{i=1}^{m}$. Then we have
@@
$\overline{\partial}_{x}\mathcal{T}x^{2}=4x_{0}\left(\frac{1-m}{2}-\sum_{i=1}^{m}\kappa(e_{i})\right)=0.$